OpenStudy (anonymous):

MEDALS!!!!!!! What is the slope of the line that passes through the points (3, –4) and (–7, 1)? a.-2 b.1/2 c.-1/2 d.2

OpenStudy (anonymous):

\[ slope=\frac{y_2-y_1} {x_2-x_1} \]

OpenStudy (anonymous):

Can you finish it?

OpenStudy (anonymous):

(3, –4) and (–7, 1)? \[ slope=\frac{1-(-4)} {-7-3}= \frac {5}{-10}=- \frac 12 \]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw